OpenStudy (anonymous):

25(x-6)=15x sovle for x

OpenStudy (anonymous):

before I give you any answers, have you atleast tied?

OpenStudy (anonymous):

first multiply out the 25(x-6)

OpenStudy (anonymous):

then put all like terms on one side

OpenStudy (anonymous):

yeah it was in my homework and i keep getting stuck and i really dont know what to do

OpenStudy (anonymous):

okay atleast you have tried

OpenStudy (anonymous):

25(x-6) = 15x 25x-150=15x 25x-15x-150=0 then put all like terms on one side by subtracting 15x from both sides then add 150 to both sides so it should look like this 25x-15x=150 than simplify the left side to 10x=150 divide both sides by 10 and you should get the value of x

OpenStudy (anonymous):

did you find what x is?

OpenStudy (anonymous):

15?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks bruh

OpenStudy (anonymous):

no prob